# Engineering subgap states in superconductors by altermagnetism

### v3

@@ Abstract
which shows that SABSs acquire a curved dispersion with $\theta$ but are not split in the presence of $d_{xy}$-AM because here $\Gamma_{\uparrow}^{+}\Gamma_{\uparrow}^{-}=\Gamma_{\downarrow}^{+}\Gamma_{\downarrow}^{-}$. Moreover, Eq. (10) further unveils that the observation of such curved SABSs require high angles, $|\theta|>\arcsin^{-1}[\Delta_{0}/\left(2h_{0}\right)]$, as demonstrated by the angle-resolved charge conductance in Fig. 4(c). The corresponding total charge conductance then exhibits a broadened ZBCP for $|h_{0}|<\Delta_{0}$ which then splits around $eV=0$ for $|h_{0}|>\Delta_{0}$, as shown in Fig. 4(d).
Figure 5: (a-c) Conductance for a junction having a chiral $d$-wave SC with a $d_{x^{2}-y^{2}}$-wave AM, including the angle-resolved charge conductance at $Z=2$, $h_{0}=\Delta_{0}$ (a), the total charge conductance as a function of eVeV for distinct $h_{0}$ (b), and the zero-bias total charge conductance as a function of $Z$ and $h_{0}$. (d-f) The same quantities as in (a-c) but with a $d_{xy}$-AM
Surface flat bands by $d_{x^{2}-y^{2}}$-wave altermagnetism in chiral $d$-wave SCs.— To have a complete picture of altermagnetism inducing subgap states, here we address it when chiral $d$-wave pairing coexists with the two types of AMs. In the absence of AMs, it is well-known that chiral $d$-wave SCs host 2 chiral SABSs [109, 110, 111, 112]. Thus, under $d_{x^{2}-y^{2}}\,$-AM, it is reasonable to expect that such SABSs are split into 4 by $d_{x^{2}-y^{2}}\,$-AM. Actually, the obtained SABSs are given by
@@
where $\nu=\pm$ and $\text{sgn}(.)$ denotes the sign operation. The split SABSs can be seen from Eq. (11). Interestingly, we verified that the dispersion of one SABS becomes steeper, while the other flattens by increasing $h_{0}$ and gets completely flat at zero energy when $h_{0}=\Delta_{0}$ for all $k_{y}$, as shown in the angle-resolved charge conductance in Fig. 5(a). The zero-energy flat band further leads to a ZBCP in the total charge conductance at $h_{0}=\Delta_{0}$ [Fig. 5(b)], while the subgap conductance becomes less prominent for $h_{0}\neq\Delta_{0}$ Fig. 5(b,c). For the $d_{xy}$-AM, the pair of SABSs acquire instead a curved dispersion due to altermagnetism which gives rise to a small enhancement of the subgap conductance, as seen in Figs.5(d,e). Moreover, such SABSs do not flatten as the $d_{xy}$-AM field enhances, which implies that the zero-biased conductance reduces by increasing $Z$ [Fig. 5(f)].
@@
Summary.— In this work, we have demonstrated rich properties of subgap bands and surface bound states where superconductivity and altermagnetism coexist. It shows that the zero-energy states can be enhanced when the symmetry of altermagnetism and pairings are the same, leading to a zero-biased conductance peak without any formation of surface bound states. We verified the robustness of the flat surface Andreev bound states arising from unconventional nodal pairings, which can be seriously influenced by either the symmetry or the magnitude of altermagnetism. It is also found that in chiral superconductors, the chiral Andreev bound states can be split, curved, or flattened by altermagnetism. In terms of experimental test of our predictions, it might be possible exploit the reported altermagnets, such as KV2Se2 [113] and various unconventional superconductors, such as cuprates [114]. Moreover, recent studies have shown that both the magnetic strength and orientation of altermagnets are electrically controllable [115, 116], providing an accessible way to change altermagnetic fields. For future applications, the subgap states are shown to bear large spin conductance with zero net magnetic field, useful for making functional superconducting spintronic devices.
@@
Acknowledgements.— B. L. acknowledges financial support from the National Natural Science Foundation of China (project 12474049). P.M. and P.B. acknowledge support by the Spanish CM “Talento Program” project No. 2019-T1/IND-14088 and No. 2023-5A/IND-28927, the Agencia Estatal de Investigación project No. PID2020-117992GA-I00 and No. CNS2022-135950 and through the “María de Maeztu” Programme for Units of Excellence in R&D (CEX2023-001316-M). J. C. acknowledges financial support from the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121) and the Carl Trygger’s Foundation (Grant No. 22: 2093). Y. T.  acknowledges financial support from JSPS with Grants-in-Aid for Scientific Research (KAKENHI Grants Nos. 23K17668, 24K00583, 24K00556, 24K00578, 25H00609 and 25H00613). Y. F. acknowledges financial support from the Sumitomo Foundation and numerical calculation support from Okayama University.

### v4

@@ Abstract
which shows that SABSs acquire a curved dispersion with $\theta$ but are not split in the presence of $d_{xy}$-wave AM because here $\Gamma_{\uparrow}^{+}\Gamma_{\uparrow}^{-}=\Gamma_{\downarrow}^{+}\Gamma_{\downarrow}^{-}$. Moreover, Eq. (10) further unveils that the observation of such curved SABSs require $|\theta|<\arcsin^{-1}[\Delta_{0}/\left(2h_{0}\right)]$, as demonstrated by the angle-resolved charge conductance in Fig. 3(b). This curved dispersion of SABSs can result in a quasi-flat band at large angle and the corresponding total charge conductance then exhibits a broadened ZBCP for $|h_{0}|<\Delta_{0}$ which then splits around $eV=0$ for $|h_{0}|>\Delta_{0}$, as shown in Fig. 3(d).
Figure 4: Conductance for junctions with chiral $d$-wave SCs with a $d$-wave AM. (a) angle-resolved charge conductance at $h_{0}=\Delta_{0}$ and (b) total conductance for distinct values of $h_{0}$ for a $d_{x^{2}-y^{2}}$-wave AM. (c) zero bias conductance vs $\alpha$ and $h_{0}$ where $\alpha$ denotes the crystal orientation of AM. (d) Same junction with (c) but the total conductance as a function of $eV$ and $\alpha$ at $h_{0}=\Delta_{0}$. $Z=2$ for all panels.
Surface flat bands by $d_{x^{2}-y^{2}}$-wave altermagnetism in chiral $d$-wave SCs.— To have a complete picture of altermagnetism inducing subgap states, here we address it when chiral $d$-wave pairing coexists with $d$-wave AMs. In the absence of AMs, it is well-known that chiral $d$-wave SCs host 2 chiral SABSs [106, 107, 108, 109]. Thus, under $d_{x^{2}-y^{2}}\,$-AM, it is reasonable to expect that such SABSs are split into 4 by $d_{x^{2}-y^{2}}\,$-AM. Actually, the obtained SABSs are given by
@@
where $\nu=\pm$ and $\text{sgn}(.)$ denotes the sign operation. The split SABSs can be seen from Eq. (11). Interestingly, we verified that the dispersion of one SABS becomes steeper, while the other flattens by increasing $h_{0}$ and gets completely flat at zero energy when $h_{0}=\Delta_{0}$ for all $k_{y}$, as shown in the angle-resolved charge conductance in Fig. 4(a). The zero-energy flat band further leads to a ZBCP in the total charge conductance at $h_{0}=\Delta_{0}$ [Fig. 4(b)], while the subgap conductance becomes less prominent for $h_{0}\neq\Delta_{0}$. For the $d_{xy}$-AM, the pair of SABSs have no spin-splittings but acquire instead a curved dispersion due to altermagnetism. It is noted that the emergence of ZBCP can be robust against a certain twist between AM and SC. As Figs.4(c,d) show, there is a conductance peak at $h_{0}=\Delta_{0}$ for a large range around $\alpha=\pi/2$ ($d_{x^{2}-y^{2}}$-AM) and disapear at $\alpha=3\pi/4$ ($d_{xy}$-AM). We thus view it as the resilient feature for experimental detection of the flat SABSs in chiral $d$-wave SCs.
@@
Summary.— In this work, we have demonstrated rich properties of subgap bands and surface bound states where superconductivity and altermagnetism coexist. It shows that the zero-energy states can be enhanced when the symmetry of altermagnetism and pairings are the same, leading to a zero-biased conductance peak without any formation of surface bound states. We verified the robustness of the flat surface Andreev bound states arising from unconventional nodal pairings, which can be seriously influenced by either the symmetry or the magnitude of altermagnetism. It is also found that in chiral superconductors, the chiral Andreev bound states can be split, curved, or flattened by altermagnetism. Our study reveals the importance of the anisotropic orbital symmetry as well as the lifting of spin degeneracy by altermagnetism in the formation of subgap states. Notably, the integration of altermagnet-superconductor heterostructures is expected to induce spin-triplet pairings [110] and the characteristics of resulting subgap states by spin-orbit and impurity effect would be an interesting topic for future research. In terms of experimental test of our predictions, it might be possible to exploit the reported altermagnets, such as KV2Se2 [97] and various unconventional superconductors, such as cuprates [111]. Moreover, recent studies have shown that both the magnetic strength and orientation of altermagnets are electrically controllable [112, 113], providing an accessible way to change altermagnetic fields. It has also been proposed that chemical doping has the potential to control the AM field [114]. For future applications, the subgap states are shown to modulate conductance spectra in the absence of a net magnetic field, useful for making field-free functional superconducting devices.
@@
Acknowledgements.— B. L. thanks F. M. Qu for useful discussions related to this work and acknowledges financial support from the National Natural Science Foundation of China (project 12474049) and Beijing National Laboratory for Condensed Matter Physics (2025BNLCMPKF011). P.M. and P.B. acknowledge support by the Spanish CM “Talento Program” project No. 2019-T1/IND-14088 and No. 2023-5A/IND-28927, the Agencia Estatal de Investigación project No. PID2020-117992GA-I00, No. PID2024-157821NB-I00 and No. CNS2022-135950 and through the “María de Maeztu” Programme for Units of Excellence in R&D (CEX2023-001316-M). J. C. acknowledges financial support from the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121) and the Carl Trygger’s Foundation (Grant No. 22: 2093). Y. T. acknowledges financial support from JSPS with Grants-in-Aid for Scientific Research (KAKENHI Grants Nos. 23K17668, 24K00583, 24K00556, 24K00578, 25H00609 and 25H00613). Y. F. acknowledges financial support from the Sumitomo Foundation and numerical calculation support from Okayama University.